HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

LINE SHAPE PARAMETERS FOR N_{2}O DILUTED IN N_{2} MEASURED BY MID-INFRARED DUAL-COMB SPECTROSCOPY

\underline{B. VISPOEL} , J. CLÉMENT, M. LEPÈRE , Research unit Lasers and Spectroscopies (LLS), Institute of Life, Earth and Environment (ILEE), University of

Namur, 61, Rue de Bruxelles, Namur, Belgium

Nitrous oxide is an important trace gas natrually present in Earth’s atmosphere but also results of human activities^{1}. It has an important role in atmospheric chemistry and in radiative transfer. N_{2}O is a strong greenhouse gas with a global warming potential 300 times higher than carbon dioxide^{2}, and contributes to the depletion of stratospheric ozone^{3}. The monitoring of its concentration requires accurate knowledge of N_{2}O spectroscopic parameters for line shape parameters can be a limiting factor of the precision of atmospheric spectrum analysis. In addition, a limited number of accurate experimental line shape parameters are needed to determine and validate intermolecular potentials used in the state-of-the-art models allowing the calculations of these parameters.

In this preliminary work, the measurements were performed using a mid-infrared high-resolution dual comb spectrometer that is based on quantum cascade lasers. It was shown that this spectrometer is well suited for line shape parameter studies^{4}. The N_{2}O-N_{2} spectra were recorded at room temperature in the 1300 cm^{−1} wavenumber region. The line shape parameters were determined by adjusting the Voigt and Speed-Dependent Voigt theoretical models on the experimental line profiles using a multi-spectrum fitting technique. The obtained results are discussed and compared to literature.

Acknowledgements: The authors thank the IRsweep company for the support during the measurements performed with the dual comb spectrometer.

1doi:10.1126/science.1179571D.J. Wuebbles, ”Nitrous Oxide: No Laughing Matter”, Science, 326(5949), 56–57 (2009)

2IPCC-Chapter 2 P. Forster et al., ”Changes in Atmospheric Constituents and in Radiative Forcing”

Chapter 2, Climate Change 2007 - The Physical Sciences Basis, Cambridge University Press, 131–217 (2007)

3doi:10.1126/science.1176985A.R. Ravishankara et al., ”Nitrous Oxide (N_{2}O): The Dominant Ozone-Depleting Substance Emitted in the 21st Century”, Science, 326, 123–125 (2009)

4doi:10.1016/j.jqsrt.2022.108239M. Lepère et al., ”A mid-infrared dual comb spectrometer in stepsweep mode for high-resolution molecular spectroscopy”, Journal of Quantitative Spectroscopy and Radiative Transfer, 287, 108239 (2022)

p-number: p229  Submitted on Wed, 21 Jun 2023 11:42:50 +0200